Kingsport Containers Company makes a single product that is subject to wide seasonal variations in demand. The company uses a job-order costing system and computes plantwide predetermined overhead rates on a quarterly basis using the number of units to be produced as the allocation base. Its estimated costs, by quarter, for the coming year are given below:

Quarter

  

First

Second

Third

Fourth

Direct materials

$

200,000

$

100,000

$

50,000

$

150,000

Direct labor

160,000

80,000

40,000

120,000

Manufacturing overhead

240,000

216,000

204,000

?

Total manufacturing costs (a)

$

600,000

$

396,000

$

294,000

$

?

Number of units to be produced (b)

160,000

80,000

40,000

120,000

Estimated unit product cost (a) ÷ (b)

$

3.75

$

4.95

$

7.35

$

?

Management finds the variation in quarterly unit product costs to be confusing and difficult to work with. It has been suggested that the problem lies with manufacturing overhead because it is the largest element of total manufacturing cost. Accordingly, you have been asked to find a more appropriate way of assigning manufacturing overhead cost to units of product.

Required:

1. Assuming the estimated variable manufacturing overhead cost per unit is $0.30, what must be the estimated total fixed manufacturing overhead cost per quarter?

2. Assuming the assumptions about cost behavior from the first three quarters hold constant, what is the estimated unit product cost for the fourth quarter?

3. What is causing the estimated unit product cost to fluctuate from one quarter to the next?

4. Assuming the company computes one predetermined overhead rate for the year rather than computing quarterly overhead rates, calculate the unit product cost for all units produced during the year.
Answer #1

Answers

  • All working forms part of the answer
  • Requirement #1

Answer: Total fixed manufacturing overhead per quarter = $ 192,000

Working

First

Second

Third

A [given]

Number of units to be produced (b)

                 160,000

                  80,000

                  40,000

B [given]

Manufacturing overhead

$              240,000

$             216,000

$            204,000

C [given]

Variable Overhead cost per unit

$                     0.30

$                    0.30

$                   0.30

D = A x C

Total variable manufacturing overhead

$                48,000

$               24,000

$               12,000

E = B - D

Fixed manufacturing overhead

$              192,000

$             192,000

$            192,000

  • Requirement #2

Answer = $ 4.15 per unit

Working

  

Fourth

A

Direct materials

$             150,000

B

Direct labor

$             120,000

C

Manufacturing overhead

$             228,000 [(120000 units x $ 0.30) + 192000]

D = A+B+C

Total manufacturing costs

$             498,000

E

Number of units to be produced

                120,000

F = D/E

Estimated unit product cost

$                    4.15 = Answer

  • Requirement #3

--The unit cost is fluctuation because:
No. of units produced are fluctuating.
As a result, Fixed Costs of $ 192000 that are fixed, gets distributed over changing no. of units.
Even though variable cost per unit remains fixed, fixed cost per unit changes with a change in no. of units produced.

  • Requirement #4

Answer: Unit Product Cost = $ 4.47

Quarter

Working

  

First

Second

Third

Fourth

TOTAL

A

Direct materials

$              200,000

$             100,000

$               50,000

$             150,000

$            500,000

B

Direct labor

$              160,000

$               80,000

$               40,000

$             120,000

$            400,000

C

Manufacturing overhead

$              240,000

$             216,000

$            204,000

$             228,000

$            888,000

D = A+B+C

Total manufacturing costs

$              600,000

$             396,000

$            294,000

$             498,000

$        1,788,000

E

Number of units to be produced

                 160,000

                  80,000

                  40,000

                120,000

              400,000

F = D/E

Estimated unit product cost

$                     3.75

$                    4.95

$                   7.35

$                    4.15

$                   4.47 = ANSWER